OpenStudy (anonymous):

Describe the solutions to y-6x^2-5x+6 A. 2 imaginary solutions B. 3 real solutions C. 1 real & 1 imaginary solution D. 2 real solutions

OpenStudy (anonymous):

6x^2 - 5x = 6 ==> 6x^2 - 5x - 6 = 0 The solutions of ax^2 + bx + c = 0 are given by: x = [-b +/- √(b^2 - 4ac)] / 2a In this case, a = 6, b = -5, and c = -6. x = [-b +/- √(b^2 - 4ac)] / 2a ==> x = {-(-5) +/- √[(-5)^2 - 4*6*(-6)]} / (2*6) ==> x = {5 +/- √[25 - (-144)]} / 12 ==> x = (5 +/- √169) / 12 ==> x = (5 +/- 13) / 12 x = (5 + 13) / 12 ==> x = 18 / 12 -> x = 1.5 x = (5 - 13) / 12 ==> x = -8 / 12 -> x = -2/3